x=y squared
(x+3)squared+y squared=53

Answer by vertciel(183)

x = y^2 ---------------(1)
(x + 3)^2 + y^2 = 53 --(2)
(1) + (2): x + (x + 3)^2 = 53
x + x^2 + 6x + 9 = 53
x^2 + 7x - 44 = 0
(x + 11)(x - 4) = 0
x = -11
x = 4
Discard x = -11 because you cannot square root a negative number to get a real number answer. Looking at (1): y^2 = -11; this will give you an imaginary number.
Therefore, y = 2 or -2 because y^2 = 4.